OpenStudy (anonymous):

(a) Express the area of a circle, A, as a function of its circumference, c. A (c)=

OpenStudy (anonymous):

Area of circle = pi*r^2 Circumference = 2*pi*r

OpenStudy (anonymous):

solve for r in the circumference equation and then plug the result into the Area of circle equation. You will end up with A (C)